OpenStudy (nightmarenight):

(fraction) Simplify: 28m^2n^4 over 8m^3n^2

OpenStudy (bibby):

Recall the property of exponents:\[\huge \frac{ x^m }{ x^n } = x^{m-n}\]

OpenStudy (nightmarenight):

so minus the exponents?

OpenStudy (bibby):

yeah basically

OpenStudy (bibby):

So apply that here. You can split the fraction up to help visually: \[\huge \frac{ 28m^2n^4 }{ 8m^3n^2}= \frac{ 7*4m^2n^4 }{ 2*4m^3n^2}\]

OpenStudy (bibby):

what cancellations do you see?

OpenStudy (nightmarenight):

m^2 and n^4 ?

OpenStudy (bibby):

yeah. also I factored out the numbers so that you can cancel the 4s

OpenStudy (nightmarenight):

is it 7n^2 over 2m

OpenStudy (bibby):

yeah good job :D

OpenStudy (nightmarenight):

Thanks for helping me! c: @bibby

OpenStudy (bibby):

anytime bby tag me next time